La procédure bonjour :

4  procedure bonjour;
     var i,z:integer;
6  begin
       for i:=1 to MAX do
8  begin
           writeln('Bonjour !');
10  z:=z+1
         end
12  writeln(z,'fois bonjour.')
     end;

Puis le corps du programme :

14 { Début du programme }
     begin
16  bonjour
     end.